What volume of wood remains if a 2-cm hole is bored out of a cube with edge lengths of 4 cm? I know to subtract the second volume from the first volume, but I can't figure out how to get the volume. Can someone please help me ASAP? I'm so confused.
Answer by rfer(12654)

v=pir^2h
v=3.14*1^2*4
v=12.56 cu cm
------------
v=s^3
v=4^3
v=64 cu cm
-------------
64-12.56=51.44 cu cm left